OpenStudy (anonymous):

You want to produce a scale drawing of your living room, which is 24 ft by 15 ft. If you use a scale of 4 in. = 6 ft, what will be the dimensions of your scale drawing? 24 in. by 10 in. 16 in. by 10 in. 24 in. by 144 in. 16 in. by 144 in.

OpenStudy (anonymous):

4 in / 6 ft = x in / 24 ft, x = 4*24/6 = 16 inches 4 in / 6 ft = y in / 15 ft, y = 15*4/6 = 10 inches